### v4

@@ 7 Demonstrations
$C_{d}=\Big{\{}\alpha:=\{x_{0},\ldots,x_{d}\},\,\,x_{i}\in\pi_{\alpha\setminus\{x_{i}\}},\,\forall i\Big{\}}.$ (25)
Proof 7.2.
@@
Proof 7.6.
It follows as a corollary from Theorem 7.3. This neighbourhood condition can be better elucidated in 7.
Theorem 7.7.
@@
Proof 7.8.
It follows as a corollary from Theorem 7.3.
Theorem 7.9.
Let $\alpha\in C_{d}$, and $N_{\alpha}$ as defined originally in Section 3. Then,
$N_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\}.$ (28)
Proof 7.10.
@@
Let $\alpha^{\prime}\in N_{\alpha}$. From Theorem 7.3, there exists $\gamma\in C_{d-1}$ such that $\gamma=\alpha\cap\alpha^{\prime}$ and in particular, $|\gamma|=d.$ Once that $|\alpha|=|\alpha^{\prime}|=d+1$ and $\gamma\subset\alpha,\alpha^{\prime}$, there exists $x\in V\setminus\alpha$ such that we can write $\alpha^{\prime}=\gamma\cup\{x\}.$ Analogously, there exists $y\in V\setminus\alpha^{\prime}$ such that we can write $\alpha=\gamma\cup\{y\}.$ It is clear that $\pi_{\gamma}\setminus(\alpha\setminus\gamma)=\pi_{\gamma}\setminus\{y\}$. Suppose this set is not empty. We need to prove that $x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma).$ By absurdity, suppose that $x$ is not in this set. By using set complementary property, we have that $x\in(\pi_{\gamma})^{\mathsf{c}}\cup(\alpha\setminus\gamma).$ If $x\in\alpha\setminus\gamma,$ then $\alpha=\alpha^{\prime}.$ If $x\in\pi_{\gamma}^{\mathsf{c}},$ then there exists $z\in\gamma$ such that $x\notin\pi_{\gamma},$ which implies that $\alpha^{\prime}$ is not a $d$-cell.
@@
The opposite inclusion is proven as follows: Let $\alpha^{\prime}$ be such that $\alpha^{\prime}=\gamma\cup\{x\},$ for some $x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma),$ and some $\gamma\in\partial(\alpha).$ Note that $|\alpha\cap\alpha^{\prime}|=|\alpha\cap(\gamma\cup\{x\})|=|(\alpha\cap\gamma)\cup(\alpha\cap\{x\})|=|\gamma|=d.$ The result follows from applying Theorem 7.3 once proven that $\alpha^{\prime}\in C_{d}.$ Indeed, $|\alpha^{\prime}|=|\gamma\cup\{x\}|=d+1.$ Suppose by absurdity that $\alpha^{\prime}\notin C_{d}.$ Than, by using the the result of Theorem 7.1, there exists $z\in\alpha^{\prime}$ such that $z\notin\pi_{\alpha^{\prime}\setminus\{z\}}.$ In case of $z=x$, we have that it contradicts that $x\in\pi_{\gamma}\setminus\{y\}$. Also, if $z\neq x,$ we would have that $z\in\gamma$ and it contradicts that $\alpha\in C_{d}.$
Theorem 7.11.
Let $\alpha\in C_{d}$, $\gamma\in\partial(\alpha)$ and $x\in\pi_{\gamma},\,x\notin\alpha$. Let $\alpha^{\prime}:=\gamma\cap\{x\}$. Then, $\alpha\in T_{\alpha}$ if and only if $x\in\pi_{\alpha}$.
Proof 7.12.
Note that Theorem 7.3 and Theorem 7.9 guarantees that $\alpha^{\prime}\in N_{\alpha}$ and that $\gamma=\alpha\cap\alpha^{\prime}$. Having said that, consider $\beta:=\alpha\cup\alpha^{\prime}=\gamma\cup\{x,y\},$ for some $y\notin\alpha,\,y\neq x$ If $x\in\pi_{\alpha},$ then $\beta\in C_{d+1},$ and $\beta>\alpha,\alpha^{\prime}.$ Otherwise, there exists $x_{0}\in\alpha$ such that $x\not\in\pi_{x_{0}}$ and then $\beta\notin C_{d+1}.$
Theorem 7.13.
Let $\alpha\in C_{d}$ and $T_{\alpha}$ as described in Section 3. Then,
$T_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\in\pi_{\alpha}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\}.$ (32)
Proof 7.14.
It follows as a corollary from Theorem 7.11 and the characterization of the neighbourhood of $\alpha$ in Theorem 7.9.
Theorem 7.15.
Let $P_{\alpha}$ as defined in Section 3. Then,
$P_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{\begin{subarray}{c}x\in\pi_{\gamma}\neq\emptyset\\ x\notin\pi_{\alpha}\neq\emptyset\\ x\notin\alpha\end{subarray}}\{\gamma\cup\{x\}\}.$ (33)
Proof 7.16.
It also follows as a corollary from Theorem 7.11 and the characterization of Theorem 7.9.
Theorem 7.17.
Let $H_{\alpha}$ as defined in Section 3. Then,
$H_{\alpha}=\bigcup_{x\in\pi_{\alpha}\neq\emptyset}\{\alpha\cup\{x\}\}.$ (34)
Proof 7.18.
@@
Let $\beta\in H_{\alpha}.$ In particular, $\beta\supset\alpha$ implies that exists $x\notin\alpha$ such that we can write $\beta=\alpha\cup\{x\}.$ It is clear that $x\in\pi_{\alpha},$ otherwise $\beta$ would not be a $(d+1)$-cell. Let $\beta$ in be a set of the form $\beta=\alpha\cup\{x\},$ for some $x\in\pi_{\alpha}.$ It is enough to show that $\beta\in C_{d+1}$, which is immediate, once that $\left|\beta\right|=d+2.$
Theorem 7.19.
@@
$\left|T_{\alpha}\right|=(d+1)\left|H_{\alpha}\right|.$ (35)
Proof 7.20.
Note that $\pi_{\alpha}=\pi_{\gamma}\cap\pi_{\alpha\setminus\gamma}$, for all $\gamma\in\partial(\alpha).$ Thus, by taking the cardinality of $T_{\alpha}$ and $H_{\alpha}$ using the equalities in Theorem 7.13 and Theorem 7.17, we have
$\left|T_{\alpha}\right|=\sum_{\gamma\in\partial(\alpha)}\left|\pi_{\alpha}\right|=(d+1)\left|\pi_{\alpha}\right|=(d+1)\left|H_{\alpha}\right|.$ (36)
Theorem 7.21.
Let $G=(V,E)$. For all $d\geq 1$ and $\alpha\in C_{d}$, the FRC is bounded by
$2\cdot(d+1)-|V|\leq\mathrm{F}_{d}(\alpha)\leq|V|$ (37)
Proof 7.22.
From (4) and (5), it is clear that we have that the minima of $F_{d}$ are reached when all neighbours of $\alpha$ are parallel. Analogously, the maxima are reached when all neighbours of $\alpha$ are transverse. Also, the maximum number of neighbours of a $d$-cell is $|V|-(d+1)$. It follows that the minima is $2\cdot(d+1)-|V|$ and the maxima is $|V|.$
Proof 7.23.
From equation (18), we know that $F_{d}(\alpha)=\frac{|T_{\alpha}|}{(d+1)}+(d+1)-|P_{\alpha}|$. Also, recall that the FRC reaches its minimum when all neighbours of $\alpha$ are parallel, i.e., when $P_{\alpha}=N_{\alpha}$, and then, $F_{d}(\alpha)=(d+1)-|N_{\alpha}|$. On the other hand, the FRC reaches its maximum when all neighbours of $\alpha$ are transverse, $T_{\alpha}=N_{\alpha}$, where in this case $|N_{\alpha}|$ is a multiple of $(d+1)$, leading us to $F_{d}(\alpha)=\frac{|N_{\alpha}|}{(d+1)}+(d+1)$. Otherwise, $|T_{\alpha}|<|N_{\alpha}|$ and $F_{d}(\alpha)\leq\left\lfloor\frac{|N_{\alpha}|}{(d+1)}\right\rfloor+(d+1)$. In the special case where $|N_{\alpha}|=|V|-(d+1)$ and $N_{\alpha}=T_{\alpha}$ (or $N_{\alpha}=P_{\alpha}$), the upper (or lower) bound is the same as in Theorem 7.21.

### v5

@@ Appendix B Demonstrations
(21) $C_{d}=\Big{\{}\alpha:=\{x_{0},\ldots,x_{d}\},\,\,x_{i}\in\pi_{\alpha\setminus\{x_{i}\}},\,\forall i\Big{\}}.$
Proof B.2.
@@
Proof B.6.
It follows as a corollary from Theorem B.3. This neighbourhood condition can be better elucidated in 7.
Theorem B.7.
@@
Proof B.8.
It follows as a corollary from Theorem B.3.
Theorem B.9.
Let $\alpha\in C_{d}$, and $N^{d}_{\alpha}$ as defined originally in Section 3. Then,
(22) $N^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma)\neq\emptyset}\{\gamma\cup\{x\}\}.$
Proof B.10.
@@
Let $\alpha^{\prime}\in N_{\alpha}^{d}$. From Theorem B.3, there exists $\gamma\in C_{d-1}$ such that $\gamma=\alpha\cap\alpha^{\prime}$ and in particular, $|\gamma|=d.$ Once that $|\alpha|=|\alpha^{\prime}|=d+1$ and $\gamma\subset\alpha,\alpha^{\prime}$, there exists $x\in V\setminus\alpha$ such that we can write $\alpha^{\prime}=\gamma\cup\{x\}.$ Analogously, there exists $y\in V\setminus\alpha^{\prime}$ such that we can write $\alpha=\gamma\cup\{y\}.$ It is clear that $\pi_{\gamma}\setminus(\alpha\setminus\gamma)=\pi_{\gamma}\setminus\{y\}$. Suppose this set is not empty. We need to prove that $x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma).$ By absurdity, suppose that $x$ is not in this set. By using set complementary property, we have that $x\in(\pi_{\gamma})^{\mathsf{c}}\cup(\alpha\setminus\gamma).$ If $x\in\alpha\setminus\gamma,$ then $\alpha=\alpha^{\prime}.$ If $x\in\pi_{\gamma}^{\mathsf{c}},$ then there exists $z\in\gamma$ such that $x\notin\pi_{\gamma},$ which implies that $\alpha^{\prime}$ is not a $d$-cell.
@@
The opposite inclusion is proven as follows: Let $\alpha^{\prime}$ be such that $\alpha^{\prime}=\gamma\cup\{x\},$ for some $x\in\pi_{\gamma}\setminus(\alpha\setminus\gamma),$ and some $\gamma\in\partial(\alpha).$ Note that $|\alpha\cap\alpha^{\prime}|=|\alpha\cap(\gamma\cup\{x\})|=|(\alpha\cap\gamma)\cup(\alpha\cap\{x\})|=|\gamma|=d.$ The result follows from applying Theorem B.3 once proven that $\alpha^{\prime}\in C_{d}.$ Indeed, $|\alpha^{\prime}|=|\gamma\cup\{x\}|=d+1.$ Suppose by absurdity that $\alpha^{\prime}\notin C_{d}.$ Than, by using the the result of Theorem B.1, there exists $z\in\alpha^{\prime}$ such that $z\notin\pi_{\alpha^{\prime}\setminus\{z\}}.$ In case of $z=x$, we have that it contradicts that $x\in\pi_{\gamma}\setminus\{y\}$. Also, if $z\neq x,$ we would have that $z\in\gamma$ and it contradicts that $\alpha\in C_{d}.$
Theorem B.11.
Let $\alpha\in C_{d}$ and $T^{d}_{\alpha}$ as described in Section 3. Then,
(23) $T^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\,\bigcup_{x\in\pi_{\gamma}\cap\pi_{\alpha\setminus\gamma}\neq\emptyset}\{\gamma\cup\{x\}\}.$
Proof B.12.
The direct inclusion is proven as follows: Let $\alpha^{\prime}\in T_{\alpha}^{d}.$ From (5), we have in particular that $\alpha^{\prime}\in N^{d}_{\alpha},$ which means by Theorem B.3 that $\gamma:=\alpha\cap\alpha^{\prime}\in C_{d-1}$ and we can characterize $\alpha^{\prime}=\gamma\cup\{x\},$ for some $x\notin\gamma$ and also $x\notin\alpha.$ We need to show that $x\in\pi_{\gamma}\cap\pi_{(\alpha\setminus\gamma)}=\pi_{\alpha}.$ If $x\notin\pi_{\alpha},$ exists $y\in\alpha$ such that $x\notin\pi_{y}$. Using that $\alpha$ and $\alpha^{\prime}$ are transverse neighbours, we have that exists $\beta\in C_{d+1}$ such that $\beta>\alpha,\alpha^{\prime},$ which means that $x,y\in\beta,$ which contradicts the fact that $\beta$ is a $(d+1)$-cell.
The opposite inclusion is proven as follows: Let $\alpha^{\prime}=\gamma\cup\{x\},$ for some $\gamma\in\partial(\alpha)$ and $x\in\pi_{\alpha}.$ We can write $\alpha=\gamma\cup\{y\},$ where $y\notin\gamma,\,y\neq x.$ Define $\beta:=\gamma\cup\{x,y\}.$ It is obvious that $\left|\beta\right|=d+2.$ We need to show that $\beta>\alpha,\alpha^{\prime}.$ Once that $\beta\supset\alpha,\alpha^{\prime},$ it is enough to show that $\beta\in C_{d+1}.$ Note that, by the definition of $\alpha^{\prime}$ and $\alpha$, $\forall z\in\gamma$ we have $z\in\pi_{\alpha}=\pi_{\beta\setminus\{y\}}$, $z\in\pi_{\alpha^{\prime}}=\pi_{\beta\setminus\{x\}}$, $x\in\pi_{\gamma}$ and $y\in\pi_{\gamma}$. In addition, since we took $x\in\pi_{\alpha}$, we have $x\in\pi_{y}$, therefore $\forall z\in\gamma\cup\{x,y\}=:\beta$, we have that $z\in\pi_{\gamma\cup\{x,y\}\setminus\{z\}}$ $=\pi_{\beta\setminus\{z\}}$.
Theorem B.13.
Let $H^{d}_{\alpha}$ as defined in Section 3. Then,
(24) $H^{d}_{\alpha}=\bigcup_{x\in\pi_{\alpha}\neq\emptyset}\{\alpha\cup\{x\}\}.$
Proof B.14.
@@
Let $\beta\in H^{d}_{\alpha}.$ In particular, $\beta\supset\alpha$ implies that exists $x\notin\alpha$ such that we can write $\beta=\alpha\cup\{x\}.$ It is clear that $x\in\pi_{\alpha},$ otherwise $\beta$ would not be a $(d+1)$-cell. Let $\beta$ in be a set of the form $\beta=\alpha\cup\{x\},$ for some $x\in\pi_{\alpha}.$ It is enough to show that $\beta\in C_{d+1}$, which is immediate, once that $\left|\beta\right|=d+2.$
Theorem B.15.
@@
(25) $\left|T^{d}_{\alpha}\right|=(d+1)\left|H^{d}_{\alpha}\right|.$
Proof B.16.
Note that $\pi_{\alpha}=\pi_{\gamma}\cap\pi_{\alpha\setminus\gamma}$, for all $\gamma\in\partial(\alpha).$ Thus, by taking the cardinality of $T^{d}_{\alpha}$ and $H^{d}_{\alpha}$ using the equalities in Theorem B.11 and Theorem B.13, we have
(26) $\left|T^{d}_{\alpha}\right|=\sum_{\gamma\in\partial(\alpha)}\left|\pi_{\alpha}\right|=(d+1)\left|\pi_{\alpha}\right|=(d+1)\left|H^{d}_{\alpha}\right|.$
Theorem B.17.
Let $P^{d}_{\alpha}$ as defined in Section 3. Then,
(27) $P^{d}_{\alpha}=\bigsqcup_{\gamma\in\partial(\alpha)}\bigcup_{\begin{subarray}{c}\Pi^{\gamma}_{\alpha}\neq\emptyset\\ x\in\Pi^{\gamma}_{\alpha}\end{subarray}}\{\gamma\cup\{x\}\},$
where
(28) $\Pi^{\gamma}_{\alpha}=\pi_{\gamma}\setminus(\pi_{\alpha}\cup(\alpha\setminus\gamma)).$
Proof B.18.
Let $\alpha^{\prime}\in C_{d}$ such that condition $1$ but not $2$ is reached. From Theorem B.3, we have that $\gamma=\alpha\cap\alpha^{\prime}\in\partial(\alpha).$ Also, we can write $\alpha^{\prime}=\gamma\cup\{x\}$ for some $x\notin\gamma$, $x\notin\alpha$ and $x\in\pi_{\gamma}$. Suppose $x\notin\Pi^{\gamma}_{\alpha}.$ Thus, using complement set property, we reach that $x\in\pi_{\alpha}\cup(\alpha\setminus\gamma)\cup(\pi_{\gamma})^{c}$. By the definition of $\alpha^{\prime}$, $x\notin\alpha\setminus\gamma$. Clearly, $x$ cannot be in $\pi_{\alpha},$ otherwise we would have $\beta=\alpha\cup\{x\}$ a $(d+1)$-cell such that $\alpha,\alpha^{\prime}<\beta$ and the second condition would be satisfied. However, if $x\notin\pi_{\gamma}$ we have that $\alpha^{\prime}\notin C_{d}.$
Now, suppose that $\alpha^{\prime}=\gamma\cup\{x\},$ some $\gamma\in\partial(\alpha)$ and $x\in\Pi^{\gamma}_{\alpha}.$ It is obvious that $\gamma<\alpha,\alpha^{\prime}.$ By absurdity, suppose that condition $2$ is reached. Then there is $\beta\in C_{d+1}$ such that $\beta>\alpha,\alpha^{\prime},$ in particular, $\beta\supset\alpha.$ However, $x\in\beta$ and $x\notin\pi_{\alpha}\cup(\alpha\setminus\gamma),$ which contradicts that $\beta$ is a $(d+1)$-cell.
Theorem B.19.
Let $\alpha=\{x_{0},\ldots,x_{d}\}\in C_{d}.$ The Forman-Ricci curvature defined in equation (6) can be re-written in terms of the node neighbourhood (1) as
(29) $F_{d}(\alpha)=(d+2)\left|\bigcap_{i=0}^{d}{}\pi_{x_{i}}\right|+2\cdot(d+1)-\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}^{d}\pi_{x_{j}}\right|.$
Proof B.20.
Using the result of Theorem B.15 and equation (5) applied to equation (7), we have that
(30) $F(\alpha)=(d+2)\left|H^{d}_{\alpha}\right|+(d+1)-\left|N^{d}_{\alpha}\right|.$
By Theorem B.9, we can have the neighbours by boundary (when they exist) in the form $\alpha^{\prime}_{i}=\gamma_{i}\cap\{y_{i}\},$ for some $y_{i}\in\pi_{\gamma_{i}}\setminus(\alpha\setminus\gamma_{i})$, for all $\gamma_{i}\in\partial(\alpha),i=1,\ldots,d.$ Once that $|\alpha\setminus\gamma_{i}|=1,\forall i=1,\ldots,d,$ we can define $\{x_{i}\}:=\alpha\setminus\gamma_{i},$ and, for all $\gamma_{i}\in\partial(\alpha),$ we have
(31) $\left|N^{d}_{\alpha}\right|=\sum_{i=0}^{d}\left|\pi_{\gamma_{i}}\setminus\{x_{i}\}\right|=\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}\pi_{x_{j}}-\{x_{i}\}\right|$
$=\sum_{i=0}^{d}\left|\bigcap_{\begin{subarray}{c}j=0\\ j\neq i\end{subarray}}\pi_{x_{j}}\right|-(d+1).$
Also,
(32) $\left|H^{d}_{\alpha}\right|=\left|\pi_{\alpha}\right|=\left|\bigcap_{i=0}^{d}\pi_{x_{i}}\right|.$
The result follows from applying these formulas to the equation (30).
Acknowledgments
This research is supported by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033. Moreover, the authors acknowledge the financial support received from the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. SR further acknowledges the RTI2018-093860-B-C21 funded by (AEI/FEDER, UE) and acronym “MathNEURO”.